CatiaDaniela25:

6) For an art show, an artist painted 153 pictures. During the show he sold 72 of his pictures. What is the ratio of pictures he still has to pictures he sold?

jhonyy9:

do you know what mean a ratio of ?

Dmoney23:

do you know the meaning of ratio

CatiaDaniela25:

no

Dmoney23:

it means a number over a number fro example : 126:189

Dmoney23:

do you get it

CatiaDaniela25:

oh ok so the answer will be 153:72

Dmoney23:

exactly

CatiaDaniela25:

oh ok thanks Dmoney23

Dmoney23:

np

Bulldogfan1234:

24

Bulldogfan1234:

you want to know what is funnir than 24

jhonyy9:

not exactly the question say what is the ratio between solded and his picture so the correct ratio will be 72 from 153 so 72/153

Bulldogfan1234:

25 ohhhhhhhhhhhh

Dmoney23:

no

Bulldogfan1234:

72:153

jhonyy9:

bc. 153 /72 not is possibile how can you sold 153 pictures from 72 ? is this possibile ?

Bulldogfan1234:

nono he painted 153 of them and sold 72 of them

Dmoney23:

no

Dmoney23:

look it up

Dmoney23:

jhonyy9

Bulldogfan1234:

no you will subtract the 72 and 153 and then get that number so it will be 81:72

Dmoney23:

you dumb

Bulldogfan1234:

me

Dmoney23:

yes

jhonyy9:

so result that he sold 72 from 153 the ratio will be 72 to 153 so 72/153 just for an example when you solve 1 from 4 of your home work problems so you solve 1/4 not 4/1

Bulldogfan1234:

no you have to subtract 153 to 72

Bulldogfan1234:

be cause you said to the pictures he still HAS to the ones he sold

loveboy:

\[\large \bf Ratio = \frac{Pictures~Left}{Pictures~sold} = \frac{153-72}{72} = \frac{81}{72}\]

loveboy:

Any doubts, do let me know :)